OpenStudy (anonymous):

(Beg. Chem.) For the reaction below, you start wtih 10 moles of C and 16 moles of D. How many moles of F do you make and what is leftover? 2C + 4D --》 3F

OpenStudy (anonymous):

The answer should be make 12 moles of F and have 2 moles of C leftover. Not sure how to get the leftover part.

OpenStudy (aaronq):

identify the limiting reactant, then subtract those moles that reacted from the reactant with more moles (not the limiting reactant).

OpenStudy (aaronq):

i know that sounds kind of confusing. the limiting reactant is D, (i divided 16/4 =4, and 10/2=5 for C) \(\dfrac{x}{2}=\dfrac{16}{4}\) x=2*4=8 so you have 10-8=2 moles left